# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v2

@@ B.3.1 Calibration Dataset
In this section, we validate the robustness of our calibration dataset design principles. We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper (Dasigi et al., 2021), MultiNews (Fabbri et al., 2019), and TREC (Li & Roth, 2002; Hovy et al., 2001). We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score. Following Section 5, all calibration datasets are constructed using the original model’s response to the context and questions as the supervision.
@@
In contrast, the RedPajama dataset without long-range dependency and model alignment shows large variance on various test sets. It also differs from the performance of the original dense model, which may incur unexpected behaviors after compression. Note that though all datasets exhibit long dependency, the questions in the TREC dataset can be answered without long context. The context in the TREC dataset of LongBench is the many-shot examples, each showing a short sentence and its classification result, while the question is to classify a new short sentence. Although the context helps to determine the complete set of 50 classes, the model can also directly clarify the sentence without any context based on common knowledge. It may contribute to a high score on the TREC test set with the RedPajama calibration dataset.
Appendix C Compression Plan Analysis
C.1 Statistics on Rules Discovered by MoA
@@
This subsection provides empirical evidence for rules discovered by MoA as mentioned in Section 6.5. The lines and spans in Figure 10 show that all heads at the first few layers generally need a high KV-Cache density. Following that, a few layers generally only require medium density. Then, in the final layers, most heads require low density, while some outlier heads need high density. This observation conforms to previous findings of the intrinsic dimension of LLM Valeriani et al. (2023). The geometry of density is similar to the intrinsic dimension of LLM, with two local minima. As observed in Figure 10, layers with lower average density (smaller values on the lines) typically display a wider range of density (wider shades). Figure 11 validates such observation. This observation confirms the need for heterogeneous attention rules within the same layer.
C.2 Connections Between MoA Rule and Semantic
In this section, we invest the masks acquired with MoA and show the interpretable semantics of the masks. Previous works manually restrict the attention pattern of the model, which may harm the semantics learned by the dense model. In contrast, MoA preserves the semantics with statistic analysis and optimization. We use visualization, human interpretation and quantitive methods to analyze the semantics of the original model and to verify whether MoA captures such semantics.
C.2.1 Mask Visualization and Semantic Categorization
Given any token, two kinds of information are used as the model inputs: position encoding and token embedding. Position encoding indicates the absolute (Zhang et al., 2022) or relative positions (Touvron et al., 2023) of tokens in the sentence. Token embedding maps different tokens as different vectors. The attention head $h$ responds to both information and output the corresponding attention value $A_{h}$. As shown in equation 6, we denote the influence of position and token of head $h$ as function $P_{h}$ and $T_{h}$, respectively. The attention value $A_{h,i,j}$ between the $i$th and $j$th token $t_{i}$ and $t_{j}$ is determined by the combination $f_{h}$ of position and token influence functions.
@@ C.2.1 Mask Visualization and Semantic Categorization
Figure 2 visualizes two typical heads that are either dominated by position $P$ or token $T$ function. For the first attention head in Figure 2, the local positional attention is clearly observed. In this head, whatever sentences are given, each token pays major attention to the first token and the prior token. As a result, the mean attention matrix accumulates extremely large attention values at the first column and the sub-diagonal. In contrast, the second attention head in Figure 2 lays more emphasis on content-based attention. Since the position distribution of important tokens are generally random, the attention matrix can show large attention values at any position. It results in a mean attention matrix without extreme mean attention values.
In conclusion, the mean attention matrix of different sentences provides a valuable insight of whether attention values of an attention head is more position-based or content-based. Intuitively, the more uneven the attention matrix value distribution is, the more position-based the head is.
C.2.2 Quantitative Semantic Analysis
We quantify how much the attention head is position-based and analyze whether MoA successfully utilizes such semantics through the evaluate-generate-optimization pipeline. We model equation 6 with a linear approximation. $P_{h}$ and $T_{h}$ are random variables with the same expectation $\mu$ and standard variance $\delta$ for all heads. For attention head $h$, the weight factor $\alpha_{h}$ evaluates the relatively influence of position and token to the final attention value.
@@ C.2.2 Quantitative Semantic Analysis
We name $\sigma_{p}(\mathbb{E}_{t}[A_{h,i,j}])$ the Standard division of Expectation (SoE) of head $h$. Note that the expectation is taken over different sentences, while the standard division is taken over different attention positions. Since $\delta_{p}$ is the same for all heads, we derive that the position impact $\alpha_{h}$ is proportional to the SoE of different heads.
The conclusion quantifies the observation stated in Section C.2.1. Intuitively, SoE shows how uneven the mean attention matrix is, thus showing the influence of position to the attention values. MoA’s generated mask density shows positive relation with SoE, suggesting that MoA successfully captures the semantic information of the dense language model as shown in Figure 12.
Appendix D Automatic Pipeline Details
D.1 Addtional Oracle on Elastic Pattern Design
@@
D.2 Derivation of Attention Influence
We use the first-order Taylor expansion to calculate the influence of each attention value. This approximation approach is supported by methodologies commonly employed in other LLM compression approaches (Li et al., 2024a; Shi et al., 2021; Das et al., 2023; Jiang et al., 2023).
@@ D.2 Derivation of Attention Influence
Given Equation 11 and  10, we derive Equation 3 as follows. For notation simplicity, we omit the head index $h$ here.
@@
D.3 Optimization Details
D.3.1 Optimizing at Single Length
@@
In this formulation, (15a) serves as the objective function to minimize the loss, subject to the constraints that each matrix selects exactly one compression plan (15b), and the average density does not exceed $d_{\mathrm{constr}}$ (15c). Finally, (15d) enforces that $X_{h,r_{h}}$ is a binary variable, indicating the selection of plans.
@@
D.3.2 Optimizing at Multiple Lengths
With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths. The key is to transform the multi-objective MIP problem into several single-objective MIP problems (Paria et al., 2018). We utilize the idea of epsilon-constraint method (Yv et al., 1971).
We first select each input length as our primary objective to perform the single-objective optimization on it while simultaneously recording the outcomes of other objectives. Specifically, for $N$ distinct objectives, we do single-objective MIP optimization on the $i$-th objective, getting minimum loss $\Delta L^{(N_{i})}_{i}$, and we concurrently collect losses of other objectives $\Delta L^{(N_{j})}_{i}$ for $j\neq i$. This process allows us to establish the range of loss $R^{(N_{j})}=\left[\min_{i}\Delta L^{(N_{j})}_{i},\max_{i}\Delta L^{(N_{j})}_{i}\right]$ for each objective. Then, we iterate through each objective again. Compared with the original multi-objective optimization in Equation 5, we now consider other objectives as constraints. To implement this, we partition each loss range $R^{(N_{j})}$ of other objectives $j\neq i$ into $M$ uniform intervals $S^{(N_{j})}_{k}$, where $0\leq k<M$. We then solve the MIP problems for each objective $i$ and iterating through the constraint intervals:
@@ D.3.2 Optimizing at Multiple Lengths
where this optimization is performed for each $i$ ranging from 00 to $N$. For each $j$, $k_{j}$ can vary independently from 00 to $M$. For efficiency consideration, we set the number of intervals as five. Finally, the results that do not conform to the Pareto front requirements are removed, resulting in the final Pareto front set of our multi-objective optimization problem.

### v3

@@ C.3.1 Calibration Dataset
Table 16: Performance comparison on various test sets, using different calibration sets. Tested on Vicuna-7B model. The result is tested with 50% density MoA on LongBench (bai2023longbench) 0-4k split.
In this section, we validate the robustness of our calibration dataset design principles. We select three sub-tasks and respective datasets from the LongBench benchmark, including Qasper (Dasigi2021Qasper), MultiNews (Fabbri2019MultiNews), and TREC (li-roth-2002-trec1; hovy-etal-2001-trec2). We use their training set to construct the calibration dataset, and use their respective test set in LongBench to calculate the score. Following Section 5, all calibration datasets are constructed using the original model’s response to the context and questions as the supervision.
@@
In contrast, the RedPajama dataset without long-range dependency and model alignment shows large variance on various test sets. It also differs from the performance of the original dense model, which may incur unexpected behaviors after the MoA search pipeline. Note that though all datasets exhibit long dependency, the questions in the TREC dataset can be answered without long context. The context in the TREC dataset of LongBench is the many-shot examples, each showing a short sentence and its classification result, while the question is to classify a new short sentence. Although the context helps to determine the complete set of 50 classes, the model can also directly clarify the sentence without any context based on common knowledge. It may contribute to a high score on the TREC test set with the RedPajama calibration dataset.
Appendix D MoA Configuration Analysis
D.1 Overview
We investigate MoA’s elastic rules for each head. As shown in Figure 11, masks in the initial and middle layers exhibit high density, aligning with conclusions from previous research on LLMs’ intrinsic dimensions (Valeriani2023hiddenRepresent) and layer sensitivities (Yuan2023ASVD). Conversely, in the final layers, most heads require low density, while few need high density. Figure 12 shows that layers with lower average density typically display more diverse densities among heads, confirming the need for heterogeneity within the same layer.
D.2 Statistics on Rules Discovered by MoA
@@
This subsection provides empirical evidence for rules discovered by MoA, as mentioned in Section D.1. The lines and spans in Figure 11 show that all heads in the first few layers generally need a high KV-Cache density and are thus assigned larger attention spans. Subsequently, a few layers generally require only medium density. In the final layers, most heads require low density, while some outlier heads need high density. This observation conforms to previous findings about the intrinsic dimension of LLMs Valeriani2023hiddenRepresent. The geometry of density is similar to the intrinsic dimension, with two local minima. As observed in Figure 11, layers with lower average density (smaller values on the lines) typically display a wider range of density (wider shades). Figure 12 validates this observation and confirms the need for heterogeneous attention rules within the same layer.
D.3 Connections Between MoA Rule and Semantic
In this section, we investigate the masks acquired with MoA and show their interpretable semantics. Previous works manually restrict the model’s attention pattern, which may harm the semantics learned by the dense model. In contrast, MoA preserves semantics with statistical analysis and optimization. We use visualization, human interpretation, and quantitative methods to analyze the semantics of the original model and verify whether MoA captures them.
D.3.1 Mask Visualization and Semantic Categorization
Given any token, two kinds of information are used as the model inputs: positional encoding and token embedding. Position encoding indicates the absolute (zhang2205opt) or relative positions (touvron2023llama2) of tokens in the sentence. Token embedding maps different tokens as different vectors. Attention head $h$ responds to both types of information and outputs the corresponding attention value $A_{h}$. As shown in equation 7, we denote the influence of position and token of head $h$ as function $P_{h}$ and $T_{h}$, respectively. The attention value $A_{h,i,j}$ between the $i$th and $j$th token $t_{i}$ and $t_{j}$ is determined by the combination $f_{h}$ of position and token influence functions.
@@ D.3.1 Mask Visualization and Semantic Categorization
Figure 2 visualizes two typical heads that are either dominated by position $P$ or token $T$ function. For the first attention head in Figure 2, local positional attention is clearly observed. In this head, regardless of the sentence, each token pays major attention to the first token and the prior token. As a result, the mean attention matrix accumulates extremely large attention values in the first column and on the sub-diagonal. In contrast, the second attention head in Figure 2 places more emphasis on content-based attention. Since the positional distribution of important tokens is generally random, the attention matrix can show large attention values at any position, resulting in a mean attention matrix without extreme values.
In conclusion, the mean attention matrix across different sentences provides valuable insight into whether an attention head is more position-based or content-based. Intuitively, the more uneven the distribution of attention matrix values is, the more position-based the head is.
D.3.2 Quantitative Semantic Analysis
We quantify how position-based an attention head is and analyze whether MoA successfully utilizes such semantics through the evaluate–generate–optimize pipeline. We model Equation 7 with a linear approximation. $P_{h}$ and $T_{h}$ are random variables with the same expectation $\mu$ and standard deviation $\delta$ for all heads. For attention head $h$, the weight factor $\alpha_{h}$ evaluates the relative influence of position and token on the final attention value.
@@ D.3.2 Quantitative Semantic Analysis
We name $\sigma_{p}(\mathbb{E}_{t}[A_{h,i,j}])$ the Standard division of Expectation (SoE) of head $h$. Note that the expectation is taken over different sentences, while the standard deviation is taken over different attention positions. Since $\delta_{p}$ is the same for all heads, we derive that the position impact $\alpha_{h}$ is proportional to the SoE of different heads.
This conclusion quantifies the observation stated in Section D.3.1. Intuitively, SoE indicates how uneven the mean attention matrix is and thus reflects the influence of position on the attention values. MoA’s generated mask density shows a positive relationship with SoE, suggesting that MoA successfully captures the semantic information of the dense language model, as shown in Figure 13.
Appendix E Automatic Pipeline Details
E.1 Additional Oracle on Elastic Pattern Design
@@
E.2 Derivation of Attention Influence
We use the first-order Taylor expansion to calculate the influence of each attention value. This approximation approach is supported by methodologies commonly employed in other LLM compression approaches (llm-mq; shi2021sparsebert; das2023beyond; jiang2023pushing).
@@ E.2 Derivation of Attention Influence
Given Equation 12 and  11, we derive Equation 3 as follows. For notation simplicity, we omit the head index $h$ here.
@@
E.3 Optimization Details
E.3.1 Optimizing at Single Length
@@
In this formulation, (16a) serves as the objective function to minimize the loss, subject to the constraints that each matrix selects exactly one MoA configuration (16b), and the average density does not exceed $d_{\mathrm{constr}}$ (16c). Finally, (16d) enforces that $X_{h,r_{h}}$ is a binary variable, indicating the selection of plans.
@@
E.3.2 Optimizing at Multiple Lengths
Figure 15: Illustration of our multi-objective Mixed-Integer Programming (MIP) approach, using a two-objective optimization example: (a) MoA first minimizes the loss for 4k inputs and records the corresponding loss for the current optimal plan at 8k. (b) Next, it minimizes the loss for 8k inputs and records the corresponding loss for the current optimal plan at 4k. These steps establish the loss ranges $R$ for both 4k and 8k input lengths. (c) MoA then re-optimizes the loss at 4k, this time using the loss intervals at 8k as different constraints. All plans generated under these constraints are recorded. (d) The last process (c) is repeated for 8k, using 4k intervals as constraints. Finally, plans meeting the Pareto front criteria for both 4k and 8k inputs are selected as the final outputs.
With the ability to optimize at a single length, we utilize the same framework for multi-objective MIP across various lengths. The key is to transform the multi-objective MIP problem into several single-objective MIP problems (Paria2018AFF). We utilize the idea of epsilon-constraint method (haimes1971bicriterion).
Figure 15 illustrates the optimization process for two input lengths. We discuss the generalized approach to handle an arbitrary number of lengths. We first select each input length as our primary objective to perform the single-objective optimization on it while simultaneously recording the outcomes of other objectives. Specifically, for $N$ distinct objectives, we do single-objective MIP optimization on the $i$-th objective, getting minimum loss $\Delta L^{(N_{i})}_{i}$, and we concurrently collect losses of other objectives $\Delta L^{(N_{j})}_{i}$ for $j\neq i$. This process allows us to establish the range of loss $R^{(N_{j})}=\left[\min_{i}\Delta L^{(N_{j})}_{i},\max_{i}\Delta L^{(N_{j})}_{i}\right$] for each objective. Then, we iterate through each objective again. Compared with the original multi-objective optimization in Equation 5, we now consider other objectives as constraints. To implement this, we partition each loss range $R^{(N_{j})}$ of other objectives $j\neq i$ into $M$ uniform intervals $S^{(N_{j})}_{k}$, where $0\leq k<M$. We then solve the MIP problems for each objective $i$ and iterating through the constraint intervals:
@@ E.3.2 Optimizing at Multiple Lengths
where this optimization is performed for each $i$ ranging from 0 to $N$. For each $j$, $k_{j}$ can vary independently from 0 to $M$. For efficiency consideration, we set the number of intervals as five. Finally, the results that do not conform to the Pareto front requirements are removed, resulting in the final Pareto front set of our multi-objective optimization problem.
E.4 Coding Interface
MoA offers a straightforward and intuitive interface for configuration search and inference. For supported HuggingFace LLM architectures, the entire search pipeline for any given new model is fully automated with a single command:
python scripts/pipeline/main.py –model_path <HF_MODEL_PATH> –model_name <NAME>
Once search is completed, the MoA configuration is as straightforward to deploy as its dense counterpart. MoA seamlessly integrates with the HuggingFace framework and remains fully compatible with high-level methods like pipeline and generate. Switching from standard attention to our mixture-structured sliding-window attention requires minimal adjustments. We provide a specialized CUDA GPU kernel for heterogeneous sliding-window attention, integrated into HuggingFace’s AttentionModule interface. The adoption involves just two lines of code:
model = update_model_function(model, model_name)
model.model.set_mixture_of_attention(moa_config)
By automating the MoA configuration search process and offering a simple deployment interface, our approach minimizes complexity and ensures practical usability.
Appendix F Limitations and Future Work
Under an extremely low-density budget, MoA fails to maintain good performance. Designing a dynamic MoA method has the potential to address this issue, which we leave for future work. Using non-linear elastic rules with bounded attention spans is also worth exploring. Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as quantization.